# BAGELS: Benchmarking the Automated Generation and Extraction of Limitations from Scholarly Text

### v1

@@ 3.1 Dataset of Extracted Limitations
Dataset Applications. The resulting dataset will be made publicly available and can be used as a benchmark for evaluating automated limitation extraction and generation methods (§5). Beyond this, the extracted limitations can be examined and organized into a taxonomy of limitations in ML and NLP, offering a more structured understanding of common research challenges. By integrating this taxonomy into citation networks, we can introduce the concept of a Limitation Multigraph, enabling scientometric analyses into whether certain limitations shape the direction of subsequent research or, alternatively, tend to be overlooked. These avenues present new opportunities to study how the reporting (or lack thereof) of limitations affects the broader scientific discourse, a topic we plan to explore in future work.
@@
4 Limitation Generation
We generated limitations by feeding every paper’s text, excluding its actual “Limitations” section into both vanilla and RAG-augmented LLMs, experimenting with either the top three sections (Abstract, Introduction, Conclusion) or the full paper as input and using three ground-truth definitions (author-stated limitations, OpenReview comments, or their combination). We computed each section’s cosine similarity to the “Limitations” section across ACL, NeurIPS dataset and found that Abstract, Introduction, and Conclusion rank highest on average (Table 19, Appendix). When full paper inputs exceed the context length, we make them chunkwise.
4.1 LLM and RAG for Limitation Generation
Most research papers do not explicitly mention limitations, or they underreport these shortcomings even when a dedicated section is provided. To address this gap, we propose a system that generates limitations directly from the paper text.
Vanilla LLM. In the vanilla LLM setup, the language model is prompted with the entire paper $P_{i}$. When the paper exceeds the context window, it is divided into chunks $\{P_{i}^{\prime}\}$, and limitations are generated for each chunk D’Arcy et al. (2024). The LLM then aggregates these chunk-specific outputs into a cohesive, meaningful final set of limitations, ensuring that no relevant material is overlooked in especially long or detailed papers.
RAG Integration. A paper $P_{i}$ can be used independently to generate limitations, but this approach risks overlooking valuable insights from other, potentially related papers. In particular, even when a paper lacks an explicit limitations section, other papers with similar methodologies or datasets may discuss relevant shortcomings. More importantly, certain findings may be implicitly contradicted by subsequent research. Also, failing to consult the broader literature leads to a significant gap in the reported limitations. To address this issue, we employ a RAG framework, which allows the system to draw on context from multiple papers rather than relying only on $P_{i}$. Specifically, we begin by splitting 100 random papers into multiple chunks and inserting these chunks into a vector database. A retriever then identifies the top $k$ chunks most relevant to the input text $\{P_{i}/P_{i}^{\prime}\}$, depending on whether the paper $P_{i}$ or a chunk $P^{\prime}_{i}$ can fit into the context window, using cosine similarity to match these chunks to the query (Figure 1, step 4,5).
@@ 5 Evaluation of Generated Limitations
We introduce a framework, called PointWise (PW), for evaluating the performance of LLMs in generating limitations for scholarly papers (Figure 2).
@@ 5.1 Coverage
[…]
Hence, $|S_{i}|=x\times y$. We then use an LLM as a judge Zheng et al. (2023) to decide if a ground truth limitation $g_{ik}$ and a generated limitation $h_{il}$ are similar in content or topic:
@@
We aggregate these coverage values across all papers by taking their means:
@@
Here, $\text{TP}_{i}$ (true positives) is the total number of matched pairs; $\text{FP}_{i}$ (false positives) is the number of ground truth limitations not matched by any LLM-generated limitation; $\text{FN}_{i}$ (false negatives) is the number of LLM-generated limitations unmatched by any ground truth limitation. True negatives ($\text{TN}_{i}$) are not defined since no “negative” class of limitations is specified. If there is one ground truth limitation $g_{i1}$ matches with multiple LLM-generated limitations (and vice versa), True Positive (TP) counts as one. (Details in Appendix A.1)
@@ 5.2 Performance
After finding matched pairs $(g_{ik},h_{il})\in M_{i}$, we evaluate their quality using text based evaluation (Rouge-1, Rouge-2, BERTSCore, Cosine Similarity), keyword based evaluation, and heading based evaluation (Details in Appendix A.2).
6 Experimental Setup
We fine-tuned three sequence-to-sequence models, T5 (512-token window) Raffel et al. (2020), BART (1024 tokens), and Pegasus (1024 tokens) Zhang et al. (2020) on a 70 / 30 train–test split. All were trained for 3 epochs with a learning rate of $5\times 10^{-5}$, weight decay 0.01, 300 warmup steps, and batch sizes of 4 (train) and 8 (eval), with early stopping; inputs longer than 512 tokens were truncated. For zero-shot evaluation, we used, GPT-3.5 (16 K) and GPT-4o (128 K) at temperature 0.1, and Llama 3 8B and Mistral 7B (128 K) via Ollama. To incorporate retrieval (RAG), we built a VectorStoreIndex with llama-index  [^10] https://www.llamaindex.ai/., and OpenAI text-embedding-ada-002 embedding model for encoding the source and query documents. We took 100 random papers from the dataset and stored them in a vector database. These documents were chunked into 800-token segments with 20-token overlaps to preserve context, and the LLM temperature was set to 0.
@@ 7 Experiments and Results
7.1 Limitation Extraction Evaluation
We manually collected limitation sections from 500 ACL papers randomly (human curated). We measure the n-gram overlap and semantic similarity between human-curated limitations vs. automatically extracted ground truth limitations to evaluate our automatic limitation extraction process. As shown in Table 1, our automatically extracted limitation text achieves strong alignment with human-curated ground truth limitations on the ACL ’23 and ACL ’24 datasets, with ROUGE-1, L scores exceeding 85% and BERTScore above 95%.
Metrics ACL 24 ACL 23
ROUGE-1 86.68 85.07
ROUGE-2 83.13 81.68
ROUGE-L 86.64 85.04
BLEU 77.70 75.15
Jaccard Similarity 79.36 77.08
JS Divergence 23.58 24.88
Levenshtein 92.54 90.32
Cosine Similarity 90.05 88.70
BERTScore (F1) 95.70 95.66
Table 1: Evaluating the Alignment of Human-Collected vs. Automated Limitation Extraction.
Human Evaluation. We employ three annotators (separate from this paper’s authors) [^11] CS graduate students with research experience in NLP and AI to evaluate the quality of the extracted limitations. We choose a sample of 100 limitations, and for each, we show them the source and ask a Yes/No question, whether they thought the LLM extracted the limitation from the source, as opposed to generating it from scratch. Each annotator answer positively in $>90\%$ of cases (avg $\pm$ std= $95\pm 2.45\%$) (Table 2). The annotator pairs have weighted Kappa scores between $38-75\%$, denoting fair to substantial agreement (Table 18, Appendix). We checked Llama 3 70b as an extractor, but the extraction quality was not good. Overall, the extraction quality is high, and the annotators agree; therefore, we use the extracted limitations as ground truth in the subsequent evaluations. Furthermore, two annotators manually verified the extracted limitations from 1000 papers. They assessed whether each automatically extracted limitation (via tool + LLM) was grounded in the original paper text or OpenReview content and whether it included any irrelevant or unrelated information. Their analysis confirmed that all extracted limitations were faithfully sourced, with no instances of hallucinated, noisy, or newly generated content.
Model Role Sample U1 U2 U3
GPT 4o mini Extractor 100 92 95 98
Table 2: Evaluating LLM as an extractor role with human annotator (U).
Model R-L BS JStext CStext CGT CLLM Prec. Recall F1
T5 19.92 87.81 10.82 31.79 35.48 29.59 0.29 0.31 0.30
BART 19.43 87.67 10.68 31.91 33.71 30.10 0.30 0.31 0.31
Pegasus 20.15 87.66 10.71 33.39 29.28 25.27 0.25 0.26 0.26
Llama 3 25.66 88.30 14.69 40.4 61.38 39.04 0.39 0.50 0.44
GPT 3.5 24.24 87.08 14.65 43.12 76.62 46.65 0.47 0.67 0.55
GPT 3.5 + RAG 30.21 90.88 19.47 45.37 39.99 44.66 0.42 0.40 0.41
GPT 4o-mini 16.57 86.02 8.70 32.29 57.65 19.76 0.20 0.31 0.24
GPT 4o mini + RAG 23.17 87.33 12.99 39.29 67.13 45.67 0.57 0.45 0.51
Table 3: Results of models in “Coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), Precision, Recall, and F1-score) and “performance” metrics – Rouge-x, BLEU, BertScore (BS), Jaccard (JS) and Cosine (CS) similarity over text (ACL data).
@@ 7.2 Limitation Generation Evaluation
Coverage Evaluation. Table 3 reports results on ACL data using only author-stated limitations as ground truth. We couldn’t collect the OpenReview of ACL papers, so we used the author-mentioned limitations as ground truth. GPT-3.5 achieves the highest Coverage of Ground Truth (CGT=76.62) and Coverage of LLM-Generated Limitations (CLLM=46.65). Vanilla LLMs tend to produce overly verbose limitations. Integrating RAG yields more focused, concise outputs that cover a broader range of topics and align better with ground truth—for both GPT-3.5 and GPT-4o mini. In particular, incorporating RAG increases CLLM by 25.91 points over GPT-4o mini. For NeurIPS data (Table 4), considering RAG increase CGT, decrease CLLM, which gains F1 score (+0.06). Compared to ACL, NeurIPS, and Peerj, the ACL dataset archives the highest CGT, and NeurIPS achieves the highest CLLM (Table 11, Appendix).
Performance Evaluation. Table 3 reports our performance metrics using author-extracted limitations as ground truth. Compared to all models, GPT 3.5 + RAG shows the best results. Augmenting GPT-3.5 with RAG boosts every measure, ROUGE-L (+5.97) and BERTScore (+3.8). GPT-4o mini exhibits the same upward trend when paired with RAG. For NeurIPS data (Table 4), considering RAG increased the performance of all metrics except Cosine Similarity. Among our datasets, NeurIPS achieves the highest ROUGE, BERTScore, and BLEU largely due to the inclusion of OpenReview feedback (Table 11, Appendix). Moreover, our proposed PointWise approach achieves better BERTScore than the traditional way in every model (Table 8, 17, Appendix).
Metric GPT 4o mini GPT 4o mini + RAG
Rouge-L 17.07 18.10 (+1.03)
BERTScore 86.25 86.45 (+0.2)
JStext 9.12 9.72 (+0.6)
CStext 35.42 34.71 (-0.71)
CGT 45.69 61.76 (+16.07)
CLLM 76.27 59.18 (-17.09)
F1 score 0.54 0.60 (+0.06)
Table 4: Coverage and Performance measure using GPT 4o mini + RAG in NeurIPS data. Ground truth is Limitation + OpenReview, and input is the top 3 sections alongside the abstract.
Evaluation ‘LLM as a Judge’. We evaluated our Judge LLM model GPT 4o mini regarding PointWise evaluation approach. From each pair of ground truth and LLM generated limitation, we asked LLM whether both are same or not (Figure 3). We use the same annotators as before and choose a sample of 200 pairs, 100 positive cases where the LLM predicted a match and 100 negatives. We ask humans the same question as the model and calculate the weighted Kappa scores. As seen in (Table 9, Appendix), the human experts overwhelmingly agree with GPT, with kappa values ranging between $90$-$95\%$. The agreement between human annotators is $>92\%$, indicating a perfect agreement. We also experimented with a strong baseline of Llama-3.1 400B, which shows a poor agreement with humans.
@@ 7.3 Ablation Study
We ablate in two dimensions a) the size of the input text. b) the ground truth
a. Size of the input text: Table 5 presents results using GPT-4o mini. Moving from the top three sections to all sections as input improves every metric across all ground-truth types. The only exceptions are a few cases, such as BERTScore and LLM Coverage (CLLM) when using OpenReview as a ground truth. We also experimented ablation study with other models (Table 6), such as Mistral 7B and Llama 3 8B on NeurIPS data. Using all sections as input yields more detailed outputs, boosting LLM coverage by 8.35 points. However, it slightly reduces ground-truth coverage, CGT (-0.29), ROUGE-L (+0.64), and BERTScore (+0.36). Llama 3 8B shows the opposite trend: using all sections significantly degrades its performance. This likely reflects Llama’s limited context window, which hinders its ability to process longer inputs and generate coherent limitations.
b. Ground truth: Table 5 compares zero-shot GPT-4o mini results against three ground-truth sets: author-stated limitations, OpenReview comments, and their combination. Using both Limitations + OpenReview yields higher scores on nearly every metric compared to Limitations alone, except for CSKeyword, JSKeyword, and CGT. This demonstrates that reviewer feedback enriches the ground truth, improving overall performance, but because the combined ground truth is larger, its relative coverage (CGT) can decrease. We also experimented with Mistral 7b (Table 7, Appendix), incorporating OpenReivew increased CLLM, Cosine similarity, and F1 score but marginal drops in other metrics.
@@
Sec Auth OPR Auth + OPR
R-L 3 16.57 18.55 17.07
(+0.5)
R-L All 17.11 (
@@
0.44
(-0.01)
Table 5: GPT 4o mini results in “coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), F1-score) and “performance” (Rouge-x, BertScore (over entire text and just heading), Jaccard (JS) and Cosine (CS) similarity over text and heading) (NeurIPS data).
Metric Input (3 Sec.) Input (All Sec.)
Mistral 7B
Rouge-L 14.59 13.95 (-0.64)
BERTScore 84.59 84.23 (−0.36)
JStext 7.70 7.49 (−0.21)
CStext 30.95 31.82 (+0.87)
BLEU 0.30 0.30 (0)
CGT 38.28 37.99 (−0.29)
CLLM 22.09 30.44 (+8.35)
F1 score 0.24 0.32 (+0.08)
Llama 3 8B
Rouge-L 25.66 22.78 (−2.88)
BERTScore 88.30 87.29 (−1.01)
JStext 14.69 11.16 (−3.53)
CStext 40.40 34.80 (−5.60)
JSkeyword 20.28 20.31 (+0.03)
CGT 61.38 18.01 (−43.37)
CLLM 39.04 5.01 (−34.03)
F1 score 0.44 0.05 (−0.39)
Table 6: Ablation study comparing Mistral 7B and Llama 3 8B on NeurIPS data. We report metrics with only the top 3 sections versus all sections as input, using author-mentioned limitations as ground truth.
8 Discussion
Our results demonstrate that leveraging external OpenReview feedback alongside author-stated limitations substantially enriches the scope and quality of LLM-generated limitations, improving performance and coverage. This indicates the model not only reproduces author-reported issues but also uncovers additional shortcomings highlighted by peer reviewers. Similarly, integrating a Retrieval-Augmented Generation (RAG) module with GPT-3.5/4 improved every metric by grounding the LLM in relevant external context. The RAG-enhanced model produced more concise, targeted limitations (fewer sentences on average), trading a slight dip in recall with improved performance. Using all sections marginally improved performance then considering 3 section in most of the metrics. Moreover, our PointWise (PW) framework breaks each limitation into fine-grained points. It employs an LLM "judge" to evaluate alignment one by one, where traditional lexical metrics can’t capture depth and relevance. Finally, chunking long documents ensures no critical content is lost outside an LLM’s context window, and RAG helps focus on significant, relevant limitations outside the current paper. Moreover, models like GPT-4o mini deliver consistently strong performance with high correlation with humans as extractors and judges, whereas others (e.g., Llama) occasionally introduce noise. These findings affirm that combining external review data, targeted section selection, RAG augmentation, and a detailed PointWise evaluation yields more trustworthy and comprehensive LLM-generated limitations—and paves the way for future work in automated scientific limitations.

### v2

@@ 3.4 Dataset Applications
The resulting dataset is publicly available [^8] https://huggingface.co/datasets/IbrahimAlAzhar/limitation-generation-dataset-bagels and can be used as a benchmark for evaluating automated limitation extraction and generation methods (§5). Beyond this, the extracted limitations can be examined and organized into a taxonomy of limitations in ML and NLP, offering a more structured understanding of common research challenges. By integrating this taxonomy into citation networks, we can introduce the concept of a Limitation Multigraph, enabling scientometric analyses into whether certain limitations shape the direction of subsequent research or, alternatively, tend to be overlooked. These avenues present new opportunities to study how the reporting (or the lack thereof) of limitations affects the broader scientific discourse, a topic we plan to explore in future work.
4 Limitation Generation
Most research papers either do not explicitly mention limitations or underreport them, even when a dedicated section is provided. We compare two systems’ ability to generate limitations from research papers: (a) vanilla LLM and (b) RAG. Note that the generators don’t have access to the text from where the limitations are extracted, e.g., limitation sections of the papers, paragraphs identified as limitations, or paper reviews; otherwise, the task would be trivial. To improve computational efficiency, we use the three most important sections of a paper as input to the generators rather than the full text. The importance score is computed by the cosine similarity of a section and a reference limitation embedding (see Table 8, Appendix).
Vanilla LLM. In the vanilla LLM setup, when the input exceeds the context window, it is divided into chunks $\{P_{i}^{\prime}\}$, and limitations are generated for each chunk D’Arcy et al. (2024). The LLM is then also asked to aggregate these chunk-specific outputs into a cohesive, meaningful final set of limitations.
RAG Integration. A paper $P_{i}$ can be used independently to generate limitations, but this approach risks overlooking valuable insights from other, potentially related papers. In particular, even when a paper lacks an explicit limitations section, other papers with similar methodologies or datasets may discuss relevant shortcomings. For example, a paper can use SVM and not explicitly mention the modeling assumptions, whereas a related paper possibly will. Moreover, certain findings may be implicitly contradicted by subsequent research. To address this issue, we employ a RAG framework, which allows the system to draw context from multiple papers rather than relying only on $P_{i}$.
There can be multiple notions of relatedness; we compare between two: a) relatedness induced by the citation network of $P_{i}$, and b) textual similarity between $P_{i}$ and other papers. For the citation network, we use both the Pcited-by, i: papers citing $P_{i}$, and Pcited-in, i: papers that $P_{i}$ cites. We parse the reference section of $P_{i}$ to extract the DOI and title of each “cited in” work. The “cited by” DOI and titles are collected from the OpenAlex API  [^9] https://openalex.org/. We query the Semantic Scholar API  [^10] https://www.semanticscholar.org/product/api with $P_{i}$’s title to get the DOIs for top 5 most semantically close papers. These DOIs and titles are cross-referenced with arXiv metadata, and the full texts of the matched papers are downloaded and parsed with the Science Parse tool.
For each paper $P_{i}$, we build separate RAG indices with a) Pcited-by, i b) Pcited-in, i, c) semantically close papers, and their combinations, where papers are split into chunks by section to preserve detail. We combine the strengths of both keyword-based (BM-25) and semantic (FAISS) search by assigning a 50% weight to the scores from each retriever. We use a LLM-based reranker, where we retrieve 20 chunks, and pass these chunks to a GPT 4o-mini model along with the original input paper. The model is prompted to score the relevance of each chunk on a scale of 1 to 10. Only the chunks that receive a relevance score of 8 or higher are ultimately selected. We compare this method with the simple baseline of just using the retrieved chunks in §7.4.
@@ 5 Evaluation of Generated Limitations
We want to evaluate the quality of the generated limitations by comparing them with the extracted ones. Functionally, both the ground-truth and the predictions are a set of text segments. NLP metrics like BERTScore, ROUGE, and cosine similarity can yield surface overlaps, providing high scores even when the generated limitations are not appropriate, too generic, or imprecise. A possible alternative is to use a holistic LLM-as-Judge approach, where the generated/ground-truth limitations are merged into single text blocks and then compared. This lacks the point-level granularity needed for fine-grained analysis. We address both these problems by introducing the PointWise (PW) evaluation framework (Figure 2).
@@ 5.1 Coverage
$S_{i}=\{(g_{ik},\,h_{il})\mid 1\leq k\leq x,\,1\leq l\leq y\}.$
Hence, $|S_{i}|=x\times y$. We then use an LLM as a judge Zheng et al. (2023) to decide if a ground truth limitation $g_{ik}$ and a generated limitation $h_{il}$ are similar in content or topic (Figure 5, Appendix):
@@
We aggregate these coverage values across all pa- pers by taking their means:
@@
Here, $\text{TP}_{i}$ (true positives) is the total number of matched pairs; $\text{FP}_{i}$ (false positives) is the number of ground truth limitations not matched by any LLM-generated limitation; $\text{FN}_{i}$ (false negatives) is the number of LLM-generated limitations unmatched by any ground truth limitation. True negative $(TN_{i})$ is not applicable in this case, as we do not have a defined negative class. If there is one ground truth limitation $g_{ik}$ that matches with multiple LLM-generated limitations (and vice versa), True Positive (TP) counts as one. (Details in Appendix A.1)
@@ 5.2 Performance
After identifying matched pairs $(g_{ik},h_{il})\in M_{i}$, we score each pair’s quality using (i) text-based metrics: ROUGE-L, BERTScore, and cosine similarity, and (ii) keyword overlap (Jaccard Similarity). Finally, the per-pair scores are averaged. Unmatched items are excluded, as our goal is to quantify similarity within aligned pairs rather than coverage (details in Appendix A.2).
6 Experimental Setup for Generation
We use three LLMs (GPT-3.5, GPT-4o-mini, and Llama 3.1 8B  [^11] https://huggingface.co/meta-llama/Llama-3.1-8B) in a zero-shot setup for both the vanilla generation and RAG. The GPT models are accessed through APIs, and the LLama models are locally deployed with Ollama. For the vanilla generation, we also fine-tune three sequence-to-sequence models, T5 (512-token window) Raffel et al. (2020), BART (1024 tokens), and Pegasus (1024 tokens) Zhang et al. (2020) on a 70 / 30 train–test split. All models were trained for 3 epochs with a learning rate of $5\times 10^{-5}$, weight decay 0.01, 300 warmup steps, and batch sizes of 4 (train) and 8 (eval), with early stopping; inputs longer than 512 tokens were truncated. For RAG, the vector database is built with llama-index  [^12] https://www.llamaindex.ai/, and the OpenAI text-embedding-ada-002 embedding model for encoding the source and query documents.
@@ 7 Experiments and Results
Evaluation of LLM as Aligner. The PointWise evaluation protocol above uses an LLM to determine whether a generated limitation matches or aligns with a ground truth one. We evaluate GPT 4o-mini’s reliability in this task. A set of 100 positive (as per the model prediction) and 100 negative instances is annotated independently by three human evaluators. The human annotators have a Cohen’s $\kappa$ score of >=>= 95% (Table 7, Appendix), which shows that the task is largely unambiguous. Cohen’s $\kappa$ between human annotators and model (GPT 4o mini) prediction is 90-95%, showing exceptional agreement. In comparison, Llama-3.1 400B shows poor agreement with the human judges (76%-81%), so in subsequent evaluations, we use GPT-4o mini as the aligner. See Table 9 in the appendix, for an example alignment.
@@ 7.1 Limitation Generation Evaluation
The ground truth contains papers that have a) only self-reported limitations and b) limitations coming from both self-reports and reviews.
7.2 Author-Mentioned Limitation
We evaluate the model’s ability to generate self-reported limitations on the ACL part of the dataset, as these papers a) have explicit limitation sections, and b) do not have open-access reviews. The results are presented in Table 3.
Vanilla LLMs and fine-tuned models. Zero-shot models outperform trained models in almost all metrics, with GPT-3.5 achieving the best results in coverage metrics, and LLama 3 achieving the best in performance metrics. Surprisingly, GPT 4o-mini has a significantly worse performance than other zero-shot models. However, the performance metrics are based on n-gram overlap and embedding measures (e.g., ROUGE, BLEU, BERTScore, cosine similarity) that primarily capture surface overlap or shallow semantics and can miss factual correctness and completeness. Therefore, we prioritize coverage-based metrics, CGT, CLLM, and F1, and report NLP metrics as secondary diagnostics.
RAG. Since GPT-3.5 performs the best in the coverage metrics, we utilize it in a RAG setup, where the index consists of “cited-in” and “cited-by” papers. This improves the performance metrics, but comes at a cost of coverage metrics. To understand whether this reduction is caused by the RAG setup or the model, we include GPT 4o-mini in the same RAG setup, which shows a significant improvement in all metrics.
Model Model type R-L BS JS CS CGT CLLM Prec. Recall F1
T5 fine-tuned 19.92 87.81 10.82 31.79 35.48 29.59 0.29 0.31 0.30
BART fine-tuned 19.43 87.67 10.68 31.91 33.71 30.10 0.30 0.31 0.31
Pegasus fine-tuned 20.15 87.66 10.71 33.39 29.28 25.27 0.25 0.26 0.26
Llama 3 zero-shot 25.66 88.30 14.69 40.4 61.38 39.04 0.39 0.50 0.44
GPT-3.5 zero-shot 24.24 87.08 14.65 43.12 76.62 46.65 0.47 0.67 0.55
GPT 4o-mini zero-shot 16.57 86.02 8.70 32.29 57.65 19.76 0.20 0.31 0.24
GPT-3.5 + RAG zero-shot + RAG 30.21 90.88 19.47 45.37 39.99 44.66 0.42 0.40 0.41
GPT 4o-mini + RAG zero-shot + RAG 23.17 87.33 12.99 39.29 67.13 45.67 0.57 0.45 0.51
Table 3: Results of models in “Coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), Precision, Recall, and F1-score) and “performance” metrics – Rouge-x, BLEU, BertScore (BS), Jaccard (JS) and Cosine (CS) similarity on the ACL dataset. In all metrics, a higher score denotes a better performance.
7.3 Self-reported & Peer-review Limitation
The ground truth here consists of author-stated limitations and peer-review limitations extracted from NeurIPS papers. We hypothesize that the RAG approaches should be beneficial for this dataset, as the reviewers are more likely to point out limitations from external sources, such as cited in/by or semantically similar papers. Therefore, we use this dataset to compare different RAG approaches with GPT 4o-mini as the baseline LLM, as the previous experiments (Table 3) suggest that it has the highest propensity of improvement with RAG.
RAG Index C_GT C_LLM F1 R-L BS CS JS
Not applicable 67.34 63.81 0.65 15.30 86.66 33.43 8.69
100 Random Papers 60.31 48.87 0.52 16.39 86.88 32.34 8.99
Cited In 69.27 62.59 0.65 14.07 86.37 33.23 8.05
Cited By 68.45 64.01 0.65 14.49 86.34 34.09 8.36
Cited In + Cited By 68.84 64.87 0.67 14.35 86.35 33.94 8.28
Cited In + By + Semantically Similar 5 Papers 68.02 63.38 0.65 14.59 86.39 33.72 8.38
Table 4: Coverage evaluation of multiple types of RAG vector database settings in NeurIPS 21-22 dataset with GPT 4o-mini as the base LLM.
Table 4 presents the performances with different RAG indices, with the first row representing vanilla LLM (no RAG). When the index is built with 100 random papers, the F1 score drops (-0.13) compared to the zero-shot approach. A combination of “cited in” and “cited by” papers achieves the highest F1 score of 0.67 – an increase of 0.02 over the baseline. However, when we further add the top five semantically related papers retrieved via the Semantic Scholar API, the F1 score reduces (-0.02), indicating that including loosely related content can introduce noise and reduce overall precision.
However, the performance metrics present a somewhat different story. Semantically related sentences from 100 randomly selected papers yield the highest scores across multiple metrics, including ROUGE-L, BERTScore, BLEU, and Jaccard similarity. We believe this is due to the vector database containing diverse texts, which are not semantically or n-gram overlapping with the ground truth. This perhaps also shows the brittleness of performance metrics for evaluating the generation quality of limitations.
@@ 7.4 Ablation Study
a. Size of the input text: We investigate the effect of the length of the input text on the generator models with an ablation study (Table 5). We use a) GPT-4o mini + RAG and b) Llama-3.1-8B, as these are the best-performing systems in the RAG and vanilla LLM setups, considering the author-mentioned limitations, the review-mentioned ones, and their combinations. When using GPT-4o-mini with RAG, expanding the context from the top-3 sections to all available sections generally increased pointwise scores for the author-written ground truth: CGT (+0.94), CLLM (+0.82), and F1 (+0.02). A similar trend was observed for the reviewer-suggested and combined ground truths, with the exception of a slight dip in the CGT score for the combined (Auth + Rev) case. By contrast, most NLP-based metrics (e.g., ROUGE, BERTScore, and cosine) slightly decreased with all-section inputs. Taken together, this indicates that using only the top-3 sections is a cost-effective alternative in this setup: minor drops in pointwise metrics, small gains (or less drop) in NLP metrics, and no large performance loss overall. In Llama-3.1-8B, however, we observe the opposite trend. Moving to all sections produces a large F1 gain (+0.13) for the combined ground truth and improves most NLP-based metrics. This suggests the smaller Llama-3.1-8B benefits from the full-paper context to generate higher-quality limitations, whereas truncating to the top-3 sections leaves it under-informed.
b. Retriever Method: On the NeurIPS dataset, we evaluate our LLM re-ranker against a vanilla retriever baseline, both operating within a RAG framework with GPT-4o mini generator (Table 6). While the baseline simply retrieves the top 3 chunks using a FAISS+BM25 search, our method re-ranks the top 20 chunks, leading to substantial gains in CGT (+28.5), CLLM (+15.53), and the F1 score (+0.24).
@@
Sec Auth Rev. Auth + Rev
GPT 4o-mini + RAG
R-L 3 16.93 11.92 14.44
All 16.73 (
$\downarrow$
12.12 (
$\uparrow$
14.35 (
$\downarrow$
BS 3 87.27 86.26 86.43
All 87.15 (
$\downarrow$
86.21 (
$\downarrow$
86.35 (
$\downarrow$
CS 3 36.03 29.16 33.88
All 35.52 (
$\downarrow$
29.81 (
@@
87.06 (
$\uparrow$
CS 3 31.82 25.16 30.72
All 31.99 (
$\uparrow$
27.47 (
$\uparrow$
29.99 (
$\downarrow$
CGT 3 63.52 42.79 44.74
All 64.47 (
$\uparrow$
57.26 (
$\uparrow$
62.04 (
$\uparrow$
CLLM 3 33.32 44.67 48.93
All 27.75 (
$\downarrow$
55.89 (
$\uparrow$
58.86 (
$\uparrow$
F1 3 0.39 0.43 0.46
All 0.34 (
$\downarrow$
0.56 (
$\uparrow$
0.59 (
$\uparrow$
Table 5: Ablation study with GPT 4o-mini + RAG and Llama 3.1 8B results in “coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), F1-score) and “performance” (Rouge-x, BertScore (BS), and Cosine (CS) similarity in the NeurIPS data.
Model VD C_GT C_LLM F1
GPT 4o-mini Vanila k =3 40.34 49.34 0.43
GPT 4o-mini LLM re-ranker 68.84 64.87 0.67
Table 6: Performance between different retriever approaches in VD (Vector Database) in RAG (vanilla RAG (considering top 3 chunks) vs LLM re-ranker)
Our findings demonstrate that a multi-faceted approach, combining curated external data with targeted retrieval, significantly enhances the generation of scientific limitations. This is especially evident when we use limitations extracted from reviews in the ground truth, as the use of “cited in” and “cited by” papers in the RAG index achieves the highest F1 score. We also observe that the length of the input to the generator model has a different effect in the vanilla LLM and RAG setup. It might be beneficial to use full paper texts for smaller models, but larger models in RAG setups can perform reasonably well with the most important parts of a paper.